EVA Report – January 27th

  

Crew 220 (MAU) EVA Report 27JAN2020

EVA # 8

Author: Johannes Svensoy, MD

Purpose of EVA: Martian Medical training EVA

Start time: 1030

End time: 1230

Narrative: The morning was spent on medical lectures by Dr.Shawna and Dr.Svensoy at the MDRS Hab. Especially focusing on P-MARCH-P for first assessment of trauma patients and secondary survey including full check of patient. We also went through SAMPLE for patient history and METHANE for calling in the incident.

The crew members were eager to test their skills on the simulated patient and on our medical manikin. Realistic training was conducted in proximity of the MDRS Hab in the sunny and windy Martian terrain. Simulated exercises included a dislocated knee after fall and crushed lower limbs after rover accident. The Golden Bubble pressurized medevac device invented by one of the crew members was used for transferring the patient to safety. The Golden Bubble pressurized medevac device can be used for suit breach and for doing medical procedures out in the Martian field.

The crew members did well and expressed valuable learning points. Accidents will happen here on Mars we are glad to have the opportunity to train on these situations before real situations rise.

At the end of the EVA crew members Matt and Shawna went over to the MAU station on an Engineering EVA to secure the area because of heavy wind.

Destination: MDRS Station and proximity.

Coordinates (use UTM NAD27 CONUS): 4250050,518500

Participants: HabComm: Morgan and Marufa (MDRS). EVA Crew: Shawna, Connie, and Matt Wise.

Road(s) and routes per MDRS Map: MDRS proximity

Mode of travel: On foot.

Support Requested: No support requested during EVA.

Science Report – January 27th

  

Crew 220 MAU Science Report 27Jan2020

Sol 8

Author: Shawna Pandya, Commander

We have been testing a portable vital sign sensor for obtaining biometrics throughout this mission and I am delighted to share that after several days of troubleshooting, the sensors are back online and gathering valuable data! These sensors have previously been tested underwater, in Antarctica and by NASA back on Earth, but this is the first time they have been tested on Mars!

In addition, we are conducting an EVA biometric wellness study, wherein we assess crew biometrics just before and after EVAs. We have gathered lots of good data so far and will continue to do so.

Tomorrow we will spend the day on geological science, briefing on soil sampling methods, then going out to collect soil samples on EVA, and finally analyzing our samples post EVA.

We are very lucky in that we share our studies with MDRS crew and we anticipate a lot of good data coming out of this mission.

Sol Summary – January 27th

  

Author’s name: Connie Delisle

Mission Status: Phenominal

Sol 8 Activity Summary:

The day started with a joint briefing and breakfast between the MAU and MDRS Crew. Today’s morning activities were designed to help the Crew put the medical training on three assessment techniques: P-MARCH-P, Secondary Site Survey and SAMPLE procedure. Crew members assessed the condition of simulated victims, and for the first time, used XO Wise’s “Golden Bubble”. The point clearly hit home – that there are many different challenges in assessing, recovering and transporting a patient on Mars.

Look Ahead Plan: Sol 9 includes another first on this Mission, a Geological EVA. First, the Crew will be briefed on soil collection procedures by Crew Member and Researcher, Morgan Kainu. Second, the Crew will attend the Science Dome to partake in soil sampling research. This includes testing soil textures and other properties along with using a black light.

Anomalies in work: The weather report did not indicate the warmth of the sun and had to redo the compression/decompression sequence at MDRS to take off cold weather clothing. Also, the wind between 011:00 and 17:00 was markedly higher than previous days, with gust estimated at 25 MPH. The Crew adapted by securing their equipment and also doing two rendezvous at MAU station. Some remedial work was needed there to ensure that the equipment was secured.

Weather: Tonight, clear skies with a lowest temperature later this evening around -6°C, with a 0% chance of snow. The average highest temperature during the day is expected to reach 5°C.

Crew Physical Status: Commanders and Crew worked very hard today – both mentally and physically. However, there is a sense of satisfaction operations become more efficient and training, science and EVAs become more prominent.

EVA: One morning and one afternoon EVA was conducted, all in front of the MDRS Station.

Reports to be filed:

Commanders Report
Sol Summary
Photo Report
Operations Report
GreenHab Report
EVA Report
EVA Request
Support Requested: None Required

Commander Report – January 27th

  


Crew 220 Commander Report 27Jan2020

Author: Connie Delisle, Commander

Title: Out of this World Medical Training

Sol 8 marked a shift from knowing about medical assessment tools to using them. What better way to close the “knowing -doing” gap than by learning from a team of two international medical doctors. Johannes Svensøy from Norway and Shawna Pandya from Canada first explained and then lead the Crew through simulated medical emergencies during today’s morning and afternoon EVA’s. Crews practised P-MARCH-P for primary trauma survey followed by more advanced techniques called the “Secondary Survey” and “SAMPLE”. Crew’s really applied themselves to put learning into action, as well as continuing to improve radio and communication skills.

XO Wise keep extremely busy – from handling daily operations, cleaning floors on top of participating in the morning medical training EVAs. History was made as the Crew tested XO Wise’s invention – “the Golden Bubble” – for the very first time under extreme conditions on Mars. XO Wise remarked that, “I was super excited to see it used in the field”. This sentiment was shared by the Crew and we look forward to hearing about the prototype’s advancement.

Between EVAs, afternoon science, evening science the schedule continues to offer challenges and opportunities for personal and professional growth. I’ve put together a series of electronic folders and to facilitate sharing of the Mission information, the Crew has open access to the folders / files through a USB key. This has enabled Commander Shawna and to collaborate efficiently and effectively. I am learning a great deal from her along with XO Svensøy, XO Wise and the Crew.

I am grateful for this opportunity to learn, share and expand my appreciation for my colleagues expertise, contributions and insights.

Commander Delisle

MDRS Station

Commander Report – January 27th

  

Title: SurMAUnting Obstacles...

Author: Shawna Pandya, Commander

Once again, I am impressed with the mental aptitude and fortitude of my crew (as well as counterparts at MDRS station). To recap, we are at a skeleton crew over the next few sols as several of our crew remain in quarantine. Despite this, both crews have pulled together to work extremely hard at maintaining station operations, while furthering mission operations at the same time.

In the morning, we started with with a recap of the medical triage and assessment skills learned to date (METHANE, P-MARCH-P), and then expanded this skillbase to include a SAMPLE history and the secondary trauma survey. Next, we put these skills into practice into the field as we drilled a trauma scenario in front of the MDRS Hab over the course of two medical EVAs. These EVAs also saw the first field deployment of the Golden Bubble pressurized medevac device, and early demonstrations are both promising and gratifying.

Life on Mars, of course, is not without its challenges. The winds picked up considerably today, and we conducted two Engineering EVAs over the course of the day to fortify and secure MAU station from the elements. I need to commend XO Svensoy in particular for his diligence and perseverance in securing the crew quarters from the elements. His presence is a true asset to the station, and we are better off for having him.

To whatever adventures may come next...

Commander Pandya
Callsign: Nightowl
MAU Station

Sol Summary – January 27th

  

MAU Crew 220 Sol 8 Summary 27Jan2020

Summary Title: Skills and Drills and SiMAUlated Fractures Oh My!
Author’s name: Shawna Pandya, Commander
Mission Status: Nominally windy
Sol Activity Summary: The MAU and MDRS rendezvoused for a day of medical drilling, skilling and practical application. In the morning, the two physicians on the crew reviewed previously learned skills, and then also taught the crew about additional triage and assessment skills. These skills were then tested in simulated trauma scenarios on EVA in front of the hab. Most excitingly, these EVAs saw the first deployment of the Golden Bubble pressurized transport device in the field, which was successful, and provided many useful data points going forward. In other news, it was a windy day on Mars, necessitating two Engineering EVAs at MAU station to safely secure the base against the elements.
Look Ahead Plan: The crew looks forward to a day full of geological science and exploration tomorrow!

Weather: -6C low tonight with light winds at 12kph / 0% chance of precipitation tonight & tomorrow
Crew Physical Status: MEDucated
EVA: AM & PM medical EVAs drilling trauma scenarios in front of the MDRS Hab
Reports to be filed:
Sol Summary (sent)
Commander's Report (sent)
EVA Report
Operations Report
GreenHab Report
Photo Report
Science Report
EVA Request: Pending.
Support Requested: MAUral support please!

Operations Report – January 27th

  



Crew 220 (MAU) Operations Report 27JAN2020

SOL 8

Name of person filing report: Johannes Svensoy, MD

Non-nominal systems: N/A

Generator: Not Run

Hours run: 0

From what time last night: N/A

To what time this morning: N/A

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): n/a

Diesel Reading – N/A

Station Propane Reading- 2%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – N/A

Water Meter (units)- N/A

Water (static tank) (gal) – N/A

Static to loft Pump used – N/A

Water in GreenHab (gal): N/A

Water in Science Dome (gal): N/A

Toilet Tank Emptied – no

Deimos rover used: N/A

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – Not used

Starting Hours:

Beginning Charge:

Ending Hours:

Ending Charge:

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Not used

Starting Hours:

Beginning Charge:

Ending Hours:

Ending Charge:

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where? Yes, post office and supplies.

CrewCar used and why, where? Not used.

General Notes and Comments: Strengthened the Hab because of strong wind.

Summary of the internet: N/A

Summary of Suit and Radios: Nominal

Summary of Hab: We have strengthened constructions because of strong winds. The floors are becoming dirty and wet due to mud and run-off. We will be reinforcing the floor over the next 2 Sols as the MAU Engineer returns from simulated quarantine. Structural integrity nominal.

Summary of Science Dome operations: Temperatures are hovering around 7°C during the day.

Summary of RAM operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: None

Operations Report – January 27th

Operations Report

Crew 220 Operations Report 27JAN2020

Sol: 8

Name of person filing report: Matt Wise

Non-nominal systems: Primary Generator, Water Heater, Hab lower deck fluorescent lights

Generator: Primary is inoperable. Rental was run-the following information is for the rental.

Hours run: 14hrs
From what time last night: 18:00

To what time this morning: 08:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 92% Not yet run

Diesel Reading – Below 1/2 (approx 48%)

Station Propane Reading- 82%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 7

Water Meter (units)- 0148360,6

Water (static tank) (gal) – 201

Static to loft Pump used – Yes

Water in GreenHab (gal): 150.34

Water in Science Dome (gal): 0

Toilet Tank Emptied – Yes

Deimos rover used: Still in the workshop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: No

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used – No

Starting Hours: 127.9

Beginning Charge: 100%

Ending Hours: 128

Ending Charge: 100%

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: No

Starting Hours: 133.9

Beginning Charge: 100%

Ending Hours: 134

Ending Charge: 100%

Notes on rovers: Opportunity and Deimos off-site for maintenance. There were no EVAs today, so the rovers were not used.

ATVs Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today: 0

Notes on ATVs: None.

Habcar used and why, where? Yes; Post Office & Supplies

CrewCar used and why, where? No

General Notes and Comments: None.

Summary of the internet: Nominal

Summary of Suit and Radios: 3 radios (black w/ blue tape) malfunctioning; one doesn’t power on, one transmits off/on rapidly when keyed, and one scrambles/screen flashes when keyed. Team conducted basic troubleshooting, power cycle, and reset settings with no effect.
Suits Nominal

Summary of Hab: Main fluorescent lights in Hab lower deck dim and flickering. Center light bar usually lighting correctly, light near door flickers and dim all times, light bar near stairs has no bulbs. Engineer attempted to remove/replace bulbs with no effect. We’re not aware of any replacement bulbs in Hab or RAM. Any further troubleshooting must be left to qualified professional

Summary of Science Dome operations: Nominal

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and Safety issues: MDRS crew all healthy

Questions, concerns, and requests to Mission Support: None. Thanks!

EVA Report – January 27th

  


Crew 220 EVA Report 27JAN2020

EVA # 8

Author: Johannes Svensoy, MD

Purpose of EVA: Martian Medical training EVA

Start time: 1400

End time: 1600

Narrative: The morning was spent on medical lectures by Dr. Shawna and Dr. Svensoy at the MDRS Hab. The lectures on P-MARCH-P for first assessment of trauma patients and secondary survey that included full check of the patient. We also went through SAMPLE for patient history and METHANE for calling in the incident.

The crew members were well fed and hydrated leaving the airlock eager to test their skills on the simulated patient and on our medical manikin. Realistic training was conducted in near proximity of the MDRS Hab in the sunny and windy Martian terrain. Simulated exercises included broken ankle after fall, consciousness patient with neck injury after fall from a ladder, and crushed lower limbs after rover accident. The Golden Bubble pressurized medevac device invented by one of the crew members was used for transferring the patient to safety. The Golden Bubble pressurized medevac device can be used for suit breach and for doing medical procedures out in the Martian field.

At the end of the EVA crew members Morgan and Johannes went over to the MAU station on an Engineer EVA to secure the site because of heavy winds.

The crew members did well and expressed valuable learning points; they report being more prepared for taking care of patients if accidents would happen.

Destination: MDRS Station and proximity.

Coordinates (use UTM NAD27 CONUS): 4250050,518500

Participants: HabComm: Shawna (MDRS). EVA Crew: Johannes, Morgan, and Marufa.

Road(s) and routes per MDRS Map: MDRS proximity

Mode of travel: On foot.

Support Requested: No support requested during EVA.

Greenhab Report – January 27th

  


 Crew 220 GreenHab Report 27-JAN-2020 

GreenHab Officers: Marufa Bhuiyan and Morgan Kainu

Environmental control: Heating. GreenHab door shade cloth (40%) (30%) on.

Average temperatures: 20.5°C

08:32

Floor Unit: 17 °C

11:30

Floor Unit: 24 °C

15:45

Floor Unit: 24 °C

18:19

Floor Unit: 17 °C

Hours of supplemental light: 19:30 to 00:00

Daily water usage for crops: 12.7 gallons

Daily water usage for research and/or other purposes: NA

Water in Blue Tank 150.34 gallons

Time(s) of watering for crops: 2

08:32

19:35

Changes to crops: Snow pea plants disposed (four plants); cucumber plant (one) is dying – permission to dispose.

Narrative: Crops are continuing to do well and thriving.

Harvest: 3g cilantro, 2g lettuce, 2g basil, 10g chard spinach, 0.5g lemon balm, 0.3g chives, 21g icicle radish, 22g butter crunch, 5g sage

Support/supplies needed: It would be awesome to have a small disposal unit for organic waste as well as a washing sink to rinse hands, freshly picked herbs, and vegetables.